OpenStudy (anonymous):

Add. (x^5 + 7x^3 – 9) + (4x^5 – 2x^3 + 13)

OpenStudy (pfenn1):

Get rid of the parentheses and add up the like terms.

OpenStudy (anonymous):

do i add the exponents too?

OpenStudy (pfenn1):

No. In multiplication you add exponents but not in addition.

OpenStudy (anonymous):

4x^5 + 9x^3 + 22 is this right

OpenStudy (pfenn1):

Would it help if I showed the expression like this?\[(x^5 +4x^5)+ (7x^3 - 2x^3) +(-9+ 13) \]

OpenStudy (anonymous):

i still dont understand, can you show the steps in finding the answer maybe

OpenStudy (pfenn1):

What is\[1x^5+4x^5=?\]

OpenStudy (anonymous):

5x^5

OpenStudy (pfenn1):

Correct! What is\[7x^3-2x^3=?\]

OpenStudy (anonymous):

5x^3

OpenStudy (pfenn1):

Correct. When adding polynomials, you must add and subtract like expressions. What is \[-9+13=?\]